# Wannier states and spin supersolid physics in the triangular antiferromagnet K2Co(SeO3)2

### v1

@@ Abstract
We use a combination of ultra-high-resolution inelastic neutron scattering and Monte Carlo numerical simulations to study the thermodynamics and the structure of spin excitations in the spin-supersolid phase of the triangular lattice XXZ easy axis antiferromagnet K2Co(SeO3)2 and its evolution in a magnetic field. BKT transitions heralding the onset of Ising and supersolid order are detected. Above the supersolid phase the value of Wannier entropy is experimentally recovered. At low temperatures, with an experimental resolution of about 23 $\mu$eV, no discrete coherent magnon modes are resolved within a broad continuum of scattering. In addition to gapless excitations, a pseudo-Goldstone mode with a small energy gap of 0.06 meV is found. A second excitation continuum is seen at higher energy, in place of single-spin-flip excitations of the Ising model. In applied fields the continuum gradually morphs into coherent spin waves, with the Goldstone and pseudo-Goldstone sectors showing distinct evolution. The agreement between experiment and numerical simulations is excellent on the quantitative level.
@@ I Introduction
Identifying and understanding novel quantum phases of matter have been a central topic in modern condensed matter physics. The search for a supersolid, an exotic quantum state which is simultaneously a solid and a superfluid [1], is one of the most celebrated examples among such efforts. Originally proposed theoretically in solid helium [2, 3], the supersolid phase has since been studied extensively across various systems, including cold atoms [4, 5, 6], hard-core bosons [7, 8, 9], and quantum spin systems [10, 11, 12, 13, 14]. Despite significant theoretical progress, experimental identification of a supersolid phase remains a challenge.
@@
The superfluid current of a bosonic gas in the BEC state corresponds to a superfluid spin current in the pure XXZ model. However, it is important to note that the spin Hamiltonian of a real spin system always includes terms that break the continuous symmetry associated with the conservation of any component of the net magnetization. Consequently, real spin systems lack a continuity equation, resulting in short-lived spin currents and making the concept of a spin superfluid poorly defined. Nevertheless, we will continue to use the term “spin supersolid” to describe spin systems that are well approximated by U(1)-invariant models, which undergo a transition into a “solid” phase at a critical temperature $T_{c1}$, followed by a second transition into a “superfluid” phase that coexists with the solid at a lower temperature $T_{c2}$. As we will see here, a key aspect that remains robust in the spin incarnation of this phenomenon is the strong quantum character of the low-energy excitations, which are not captured by a semi-classical $1/S$ expansion.
@@
Figure 1: (a) Classical Y-phase. The order breaks the U(1) symmetry of the Hamiltonian selecting one orientation on the $xy$-plane and the ${\text{C}}_{3}$ lattice geometry selecting one sublattice for the spin down. (b) The classical ground state exhibits accidental continuous degeneracy due to the freedom to select one of the three spin angles, with the other two constrained by the derived relation Eq. 2.
While the spin supersolid ground state is well understood theoretically, the spectrum of magnetic excitations remains an open question. Recent inelastic neutron scattering (INS) study on the triangular lattice easy-axis XXZ antiferromagnet K2Co(SeO3)2 [28] sheds new light on this issue. The ground state of K2Co(SeO3)2 has been proposed as a spin supersolid phase with a “Y” structure [28, 29, 30]. Interestingly, the spin dynamics is highly unusual, with the low-energy magnetic spectrum dominated by a broad continuum of excitations, aside from the gapless Goldstone modes, and a roton-like excitation dip observed at the M points of the Brillouin zone.
@@
II Material and Method
K2Co(SeO3)2 is an easy-axis $S=1/2$ antiferromagnet very close to the Ising limit with $\alpha=0.08$ [28] deduced from high-field magnetometry measurements. At zero magnetic field, it exhibits no long-range magnetic order down to $T=$ 0.35 K [30]. Below that our calorimetric study revealed an additional low-temperature phase [28], which has been proposed to be the theoretically predicted spin-supersolid [24, 16]. When a magnetic field is applied along the easy axis, a quantum critical point is found at $\mu_{0}H_{c}=0.8(1)$ T, beyond which the system transitions into the uud plateau phase. Neutron diffraction studies have revealed quasi-two-dimensional magnetic reflections with propagation vectors (1/3, 1/3) in both the Y- and the uud phases, suggesting magnetic structures with a $\sqrt{3}\times\sqrt{3}$ supercell qualitatively in agreement with the expected three-sublattice order in the spin-supersolid and spin-solid phase [28].
@@ II Material and Method
Figure 2: A schematic view of the crystal structure of $\text{K}_{2}\text{Co}(\text{Se}\text{O}_{3})_{2}$. (a) Unit cell with ABC-stacked triangular planes of Co2+ ions. (b) Top view of a single triangular plane. Inset: photo of a $\text{K}_{2}\text{Co}(\text{Se}\text{O}_{3})_{2}$ single crystal on a millimeter grid paper.
Single crystals of K2Co(SeO3)2 are grown as in Ref. [30]. The crystal structure contains ABC-stacked two-dimensional triangular lattice planes of Co2+ ions separated by K layers [Figure 2]. The space group is $R$-3m, and the lattice parameters are $a=5.52$, and $c=18.52$ [36]. High-resolution inelastic neutron scattering measurements were performed using the AMATERAS cold-neutron time-of-flight spectrometer [37] in J-PARC, Japan. The incident neutron energies used are $E_{i}$ = $1.69$ and $7.73$ meV. The energy resolution at elastic scattering is $23$ and $240$ $\mu$eV, respectively. The data were collected by rotating the sample by $160$ degrees with $1$ degree per step, and analyzed using the Horace package [38]. Inelastic neutron scattering measurements in a magnetic field were carried out using the LET cold-neutron time-of-flight spectrometer in ISIS facility, Rutherford Appleton Laboratory, United Kingdom. The incident neutron energy is fixed at $E_{i}$ = 2.15 meV. The spectrometer is operated in the high-flux mode, with chopper frequency 240/120 Hz, yielding an energy resolution of 40 $\mu$eV at elastic scattering. The data were obtained by rotating the sample by 90 degrees with 1 deg per step. A single crystal sample of 425 mg and 1.23 grams were used for AMATERAS and LET, respectively. They are mounted on a copper holder and cooled using a 3He/4He dilution refrigerator. At LET, a magnetic field is applied along the [0,0,1] direction using a vertical-field cryomagnet. In both cases the sample was oriented with [1,0,0] and [0,1,0] directions in the horizontal plane. The specific heat measurements were performed using a relaxation method by a Quantum Design Physical Property Measurement System on a 0.05 mg single-crystal sample for 0.12--4 K with a dilution fridge insert and on a 3.95 mg one for 2--300 K.
All LSWT calculations were performed using Sunny.jl [39], worldline QMC calculations were performed using a modified directed-loop algorithm [40, 41] for system sizes of $36\times 36$ (thermodynamic properties) and $72\times 72$ (spin dynamics). The analytic continuation of the QMC data was carried out using the SmoQyDEAC.jl library [42, 43].
III Model and theory

### v2

# thanks: These two authors contributed equallythanks: These two authors contributed equallythanks:

@@ Abstract
We combine ultra-high-resolution inelastic neutron scattering and quantum Monte Carlo simulations to study thermodynamics and spin excitations in the spin-supersolid phase of the triangular lattice XXZ antiferromagnet K2Co(SeO3)2 under zero and non-zero magnetic field. BKT transitions signaling the onset of Ising and supersolid order are clearly identified, and the Wannier entropy is experimentally recovered just above the supersolid phase. At low temperatures, with an experimental resolution of about 23 $\mu$eV, no discrete coherent magnon modes are resolved within a broad scattering continuum. Alongside gapless excitations, a pseudo-Goldstone mode with a 0.06 meV gap is observed. A second, higher-energy continuum replaces single-spin-flip excitations of the Ising model. Under applied fields, the continuum evolves into coherent spin waves, with Goldstone and pseudo-Goldstone sectors responding differently. The experiments and simulations show excellent quantitative agreement.
@@ I Introduction
Identifying and understanding novel quantum phases of matter has been a central topic in modern condensed matter physics. The search for a supersolid, an exotic quantum state which is simultaneously a solid and a superfluid [1], is one of the most celebrated examples among such efforts. Originally proposed theoretically in solid helium [2, 3], the supersolid phase has since been studied extensively across various systems, including cold atoms [4, 5, 6], hard-core bosons [7, 8, 9], and quantum spin systems [10, 11, 12, 13, 14]. Despite significant theoretical progress, experimental identification of a supersolid phase remains a challenge.
@@
The superfluid current of a bosonic gas in the BEC state corresponds to a superfluid spin current in the pure XXZ model. However, it is important to note that the spin Hamiltonian of a real spin system always includes terms that break the continuous symmetry associated with the conservation of any component of the net magnetization. Consequently, real spin systems lack a continuity equation, resulting in short-lived spin currents and making the concept of a spin superfluid poorly defined. Nevertheless, we will continue to use the term “spin supersolid” to describe spin systems that are well approximated by U(1)-invariant models, which undergo a transition into a “solid” phase at a critical temperature $T_{c1}$, followed by a second transition into a “superfluid” phase that coexists with the solid at a lower temperature $T_{c2}$. As we will see here, a key aspect that remains robust in the spin incarnation of this phenomenon is the strong quantum character of the low-energy excitations.
@@
Figure 1: Classical ground state magnetic structures. (a) Classical Y-phase. The order breaks the U(1) symmetry of the Hamiltonian selecting one orientation on the $xy$-plane and the ${\text{C}}_{3}$ lattice geometry selecting one sublattice for the spin down. (b) The classical ground state exhibits accidental continuous degeneracy due to the freedom to select one of the three spin angles, with the other two constrained by the derived relation Eq. 2.
While the spin supersolid ground state is well understood theoretically [26, 16, 29, 9, 30, 31, 7, 32], the spectrum of magnetic excitations remains an open question. Recent inelastic neutron scattering (INS) study on the triangular lattice easy-axis XXZ antiferromagnet K2Co(SeO3)2 [33] sheds new light on this issue. The ground state of K2Co(SeO3)2 has been proposed as a spin supersolid phase with a “Y” structure [33, 34, 35]. Interestingly, the spin dynamics is highly unusual, with the low-energy magnetic spectrum dominated by a broad continuum of excitations, aside from the gapless Goldstone modes, and a roton-like excitation dip observed at the M points of the Brillouin zone.
@@
Figure 2: Crystal structure of K2Co(SeO3)2. A schematic view of the crystal structure of $\text{K}_{2}\text{Co}(\text{Se}\text{O}_{3})_{2}$. (a) Unit cell with ABC-stacked triangular planes of Co2+ ions. (b) Top view of a single triangular plane. Inset: photo of a $\text{K}_{2}\text{Co}(\text{Se}\text{O}_{3})_{2}$ single crystal on a millimeter grid paper.
Figure 3: Temperature dependence of specific heat in K2Co(SeO3)2. (a) Heat capacity measured in K2Co(SeO3)2 (shaded circles) and in the non-magnetic counterpart K2Mg(SeO3)2 (solid squares) as a function of temperature. A 1.08 scale factor was applied to the latter, to match the heat capacities of the two materials in the range 50 to 100 K. (b) Temperature dependence of the magnetic contribution to specific heat in K2Co(SeO3)2. The phonon contribution was estimated based on the scaled K2Mg(SeO3)2 data. The two prominent peaks correspond to the two higher-temperature BKT transitions. Inset: expanded view of the low-temperature magnetic heat capacity. A weaker feature at lower temperature marks the third transition. Actual BKT transition temperatures are typically lower than the corresponding specific heat maximum [41].
Figure 4: Magnetic entropy as a function of temperature in K2Co(SeO3)2. Temperature dependence of magnetic entropy in $\text{K}_{2}\text{Co}(\text{Se}\text{O}_{3})_{2}$ in zero applied field, obtained by integrating the measured magnetic specific heat $C_{\text{mag}}/T$. The dashed lines indicate the Wannier entropy and the expected high-temperature value $\ln{2}$ ). The inset is a zoomed-in view to highlight the plateau at the Wannier entropy. The 10% discrepancy between the high-temperature data and the expected saturation value is attributed to imperfect phonon subtraction. This effect is expected to be much smaller at temperatures below 10 K.
II Results
II.1 Material
K2Co(SeO3)2 is an easy-axis $S=1/2$ triangular lattice antiferromagnet [see Figure 2] very close to the Ising limit with $J_{zz}$ = 3.1 meV, $\alpha=0.08$ [33] deduced from high-field magnetometry measurements. At zero magnetic field, it exhibits no long-range magnetic order down to $T=$ 0.35 K [35]. Below that our calorimetric study revealed an additional low-temperature phase [33], which has been proposed to be the theoretically predicted spin-supersolid [25, 16]. When a magnetic field is applied along the easy axis, a quantum critical point is found at $\mu_{0}H_{c}=0.8(1)$ T, beyond which the system transitions into the uud plateau phase. Neutron diffraction studies have revealed quasi-two-dimensional magnetic reflections with propagation vectors (1/3, 1/3) in both the Y- and the uud phases, suggesting magnetic structures with a $\sqrt{3}\times\sqrt{3}$ supercell qualitatively in agreement with the expected three-sublattice order in the spin-supersolid and spin-solid phase [33].
II.2 Specific heat and entropy
Figure 3(a) shows the zero-field specific heat of K2Co(SeO3)2 measured up to high temperatures (open circles). The phonon contribution is approximated by the measured specific heat of the isostructural nonmagnetic counterpart K2Mg(SeO3)2 (solid squares). An empirical scale factor of 1.08 was applied in order to match the heat capacities of the two systems in the temperature range 50--100 K. The upturn below 0.2 K [42] is ascribed to a nuclear Schottky-like feature (see Supplementary Note 1). The magnetic heat capacity $C_{\text{mag}}$ of K2Co(SeO3)2 is plotted as a function of temperature in Figure 3(b). Here the phonon and the nuclear Schottky contributions at the lowest temperatures were subtracted from the total heat capacity (see Supplementary Figure 1). Signatures of three distinct phase transitions are observed at $T_{\text{BKT,1}}\approx$ 11 K, $T_{\text{BKT,2}}\approx$ 0.8 K, and $T_{\text{BKT,3}}\approx$ 0.35 K (dashed lines). While the third transition is somewhat weakly defined, its position exactly corresponds to the zero-field asymptotic of a much sharper lambda-anomaly in applied field [33]. As to be discussed in more detail in Sec. II.4, these features are BKT transitions corresponding to the Ising and XY orderings of the supersolid phase, respectively.
In Figure 4 we show the magnetic entropy $\Delta S(T)$ of K2Co(SeO3)2 as a function of temperature, obtained by integrating the magnetic heat capacity $C_{\text{mag}}/T$ from the lowest temperature measured. A plateau is observed at $T_{\rm cr}\simeq 2.5~{}\text{K}$. This crossover temperature is of similar order of magnitude as the exchange constant for the in-plane spin component $J_{xy}/k_{B}$, which suggests a separation of energy scales in the magnetic excitation spectrum. The plateau value, $\Delta S/R\approx$ 0.323, aligns well with the zero-point entropy of the pure Ising model calculated by Wannier. This implies that the low-energy excitations arise from the “Wannier states”.
Figure 5: High-resolution magnetic excitation spectrum of K2Co(SeO3)2 in the supersolid phase. False-color plot of inelastic neutron scattering intensity measured in K2Co(SeO3)2 in zero applied field and $T=0.07$ K as a function of energy and wave vector transfers along high-symmetry directions on the AMATERAS spectrometer. The incident neutron energy is $E_{i}$ = 1.69 meV. The energy resolution is 23 $\mu$eV at the elastic line. This is to be compared to the 36 $\mu$eV resolution of the experiments described in Ref. [33]. The data are integrated over ±$\pm 0.5$ r.l.u. in $l$ and ±$\pm 0.05$ r.l.u. in the $(h,k,0)$ plane perpendicular to the cut, and averaged over equivalent paths shown in Supplementary Figure 3.  No background has been subtracted. Both longitudinal and transverse spin fluctuations contribute to scattering, the intensity being proportional to $S^{zz}(\mathbf{q},\omega)+S^{\bot}(\mathbf{q},\omega)$. White solid lines are the dispersion curves calculated by linear spin wave theory. Circles are first poles of $S^{zz}(\mathbf{q},\omega)$ calculated by QMC. The white hexagon represents the Brillouin zone boundary. $\Gamma$, K and M label high symmetry points in the reciprocal space. Inset: a zoom-in view of the low-energy spectrum near the K point with a different color scale to highlight the gapped pseudo-Goldstone excitation. Labels on the top axis indicate the position of energy cuts shown in Figure 6.
II.3 Inelastic neutron scattering
[…]
The continuum of excitations is also obvious in the energy-dependent cuts plotted in Figure 6 at several representative wave vectors: M point [Figure 6(b)], half way between the K and M points [Figure 6(c)], and half way between the M and $\Gamma$ points [Figure 6(e)]. A sharp excitation feature may be present at the lower boundary of the continuum for these wave vectors, whereas it can not be resolved at the wave vector halfway between the $\Gamma$ and K points [Figure 6(d)]. Note that the intensity of the scattered neutrons is plotted in the logarithmic scale, and the corresponding plots with linear scale are displayed in the insets. The observation of a sharp feature at the lower boundary of a continuum of excitations in low-dimensional quantum magnets is often interpreted as a bound state of fractionalized spin excitations [43]. In the present study, we cannot resolve any gap between the sharp excitation and the rest of the continuum within our experimental resolution, which seems to support such a scenario. Nevertheless, one can never exclude the possibility that there exist closely spaced coherent magnon modes where the difference in energy is much smaller than our experimental resolution. Even then, our high-resolution INS data imposed strong constraints on the upper limit of the energy spacing between potential discrete magnon modes. Further, our data allow us to make an accurate estimate on the spectral weight contributions of the sharp magnon modes, if existing, at the lower boundary of the continuum. At the wave vectors investigated [Figure 6(b),(c),(e)], the weight of the sharp excitation (represented by a shaded Gaussian in which FWHM corresponds to the energy resolution) is only about 40%, implying that at least 60% of the total spectral weight is contributed by the continuum.
Figure 6: Energy cuts of the magnetic excitation spectrum in K2Co(SeO3)2 in the supersolid phase. Energy cuts through the INS data shown in Figure 5 at representative wave vectors as indicated. Both longitudinal and transverse spin fluctuations contribute to scattering, the intensity being proportional to $S^{zz}(\mathbf{q},\omega)+S^{\bot}(\mathbf{q},\omega)$. (a) K point, (b) M point, (c) half way between K and M, (d) half way between $\Gamma$ and K, and (e) half way between M and $\Gamma$. The solid lines represent QMC calculations for $S^{zz}(\mathbf{q},\omega)$ only. An arbitrary global scale factor is applied to match the intensity of the M point. The measured spectra have been averaged over equivalent wave vectors accessible in the experiment. Note the y-axis is in logarithmic scale. Insets show the same plot with y-axis in linear scale. Shaded Gaussians represent the energy resolution. The dashed lines are fits to a constant background and the elastic scattering peak.
At the K point, in addition to the gapless Goldstone excitations as expected from the spontaneously broken U(1) symmetry in the supersolid phase, a gapped excitation is observed at approximately $E\approx 0.06$ meV [Figure 6(a)]. This excitation corresponds to a pseudo-Goldstone mode and is prominently visible in the energy-momentum plot of the magnetic excitation spectrum near the K point, as illustrated in the inset of Figure 5.
Apart from these low-energy excitations, we also observe higher-energy excitations near 3 meV, using incident neutron energy $E_{i}=7.73$ meV, as shown in Figure 7. In the pure Ising model, dispersionless single-magnon bands are predicted at energies $E=nJ_{zz}$ with $n=1,2,3$, corresponding to spin-flip excitations within a locally magnetized environment as illustrated in Figure 8. The observed high-energy excitation corresponds to the spin flip excitation with $n=1$. It becomes slightly dispersive due to the nonzero $J_{xy}$ ($\alpha\approx 0.08$ in K2Co(SeO3)2), with the bottom of the branch located at the K point. The average energy of the excitation is consistent with the exchange interaction $J_{zz}=3.1$ meV, as deduced from high-field magnetization data [33]. Instead of a sharp coherent mode, this excitation manifests as a gapped continuum, with its broadening attributed to strong quantum fluctuations within the Wannier sector. Similar to the lower-energy excitations, a sharp mode might exist at the lower bound of the continuum, as suggested by the shaded Gaussians in the energy-dependent cuts plotted in Figure 9. Note that to enhance the counting statistics of the neutron intensity, the INS data presented in Fig. 5-9 have been averaged over equivalent paths in the reciprocal space accessible in the experiment, as shown in Supplementary Figure 3 and Supplementary Figure 5.
Figure 7: High-energy magnetic excitations in K2Co(SeO3)2 in the supersolid phase. False-color plot of neutron scattering intensity measured on K2Co(SeO3)2 at the AMATERAS spectrometer with $E_{i}$ = 7.73 meV at $T=70$ mK and $H$ = 0, as a function of energy and wave vector transfers along high-symmetry directions in the Brillouin zone. The energy resolution is 0.24 meV at the elastic position. The data were integrated over ±$\pm 0.5$ r.l.u. in $l$ and ±$\pm 0.05$ r.l.u. in the $(h,k,0)$ plane perpendicular to the cut, and averaged over equivalent paths shown in Supplementary Figure 5.  No background has been subtracted. White hexagon represents the Brillouin zone boundary. $\Gamma$, K and M label high symmetry points in the reciprocal space. Labels above the top axis indicate the positions of energy-cuts shown in Figure 9.
Figure 8: High energy excitations at zero field. Sketch of the high-energy excitations in the Ising model at zero field. The excitation corresponds to a flip in the central blue spin changing the net magnetization in ±$\pm 1$ and the energy in (a) $J_{zz}$; (b) $2J_{zz}$; (c) $3J_{zz}$.
Figure 9: Energy cuts of the high-energy magnetic excitations in K2Co(SeO3)2 in the supersolid phase. Energy-cuts through the INS data shown in Figure 7 at representative wave vectors: (a) K point, (b) M point, (c) half way between $\Gamma$ and K, and (d) half way between M and $\Gamma$. The wave vectors are indicated in Figure 7 by arrows. The spectra have been averaged over equivalent wave vectors. Note the y-axis is in logarithmic scale. The inset shows the same plot with $y$-axis on a linear scale. The shaded Gaussian represents the energy resolution. The dashed lines are empirical fits to the background.
Figure 10: Evolution of the magnetic excitation spectrum of K2Co(SeO3)2 in the supersolid phase in a magnetic field. False-color plots of neutron scattering intensity measured on K2Co(SeO3)2 on the LET spectrometer at $T$ = 50 mK as a function of energy and wave vector along high-symmetry directions of the Brillouin zone in a magnetic field applied along the $c$ axis: (a) $\mu_{0}H=0$, (b) 0.25 T, (c) 0.5 T, and (d) 0.75 T. The incident neutron energy is $E_{i}$ = 2.15 meV, the resolution is 0.04 meV at the elastic line. The data were integrated over ±$\pm 0.75$ r.l.u. in $l$ and ±$\pm 0.05$ r.l.u. in the $(h,k,0)$ plane perpendicular to the cut, and averaged over equivalent paths shown in Supplementary Figure 4. The spectra measured at 7 T have been subtracted to remove the background from the sample environment. Note that in (d) the residual scattering at the K point near 0.4 meV is due to spurious scattering from the cryomagnet and imperfect background subtraction. Solid lines are the dispersion curves calculated by the linear spin wave theory. Circles are first poles of $S^{zz}(\mathbf{q},\omega)$ calculated by QMC. The white hexagon represents the Brillouin zone boundary. $\Gamma$, $K$ and $M$ label high symmetry points in the reciprocal space.
Figure 11: Energy cuts of the magnetic excitation spectrum in K2Co(SeO3)2 in a magnetic field. Magnetic field evolution of the neutron scattering intensity in K2Co(SeO3)2, as measured at the LET spectrometer with $E_{i}=2.15$ meV, as a function of energy transfer at the (a-d) K point and (e-h) M point. A magnetic field of $\mu_{0}H$ = 0, 0.25, 0.5, and 0.75 T, respectively, is applied along the crystallographic $c$ axis. The spectra have been averaged over equivalent wave vectors accessible in the experiment. The spectra measured at 7 T have been subtracted to remove the background from the sample environment. Note in (d) the scattering near 0.4 meV is spurious and due to imperfect backbround subtraction. The scale on the y-axis is linear. The shaded Gaussians represent the calculated energy resolution. The solid lines represent QMC calculations for $S^{zz}(\mathbf{q},\omega)$ only. An arbitrary global scale factor is applied to match the intensity of the M point at $H$ = 0. The dashed lines are fits to the elastic peak.
To better understand the nature of the low-energy continuum of excitations in the supersolid phase of K2Co(SeO3)2, we have investigated their evolution in a magnetic field applied along the magnetic easy axis (i.e. $c$ axis) at LET. In Figure 10 (a)-(d), we show the false-color plots of the energy-momentum cuts of the low-energy magnetic excitation spectra measured at $T$ = 50 mK and in a field of $\mu_{0}H$ = 0, 0.25, 0.5, and 0.75 T, respectively. The spectra were plotted after subtracting the background scattering from the sample environment measured at $\mu_{0}H$ = 7 T (note there is no magnetic excitation below 1 meV at 7 T). As before, to enhance the counting statistics, the spectra presented have been averaged over equivalent paths in the reciprocal space as illustrated in Supplementary Figure 4. Typical averaged energy-dependent cuts of the spectra at the K and M points are shown in Figure 11. The energy resolution of this dataset ($\Delta E\approx 40$ $\mu$eV) is lower than that obtained at AMATERAS [Figure 5 and Figure 6]. While the low-energy spectrum at zero magnetic field [Figure 10(a) and Figure 11(a),(e)] appears qualitatively similar in both measurements, the application of a magnetic field drastically changes the excitation spectrum.
[…]
At $\mu_{0}H=0.5$ T, the magnetic excitation spectrum undergoes further reconstruction, as shown in Figure 10(c). The gapped excitation at the K point shifts to a higher energy of 0.26 meV [see also Figure 11(c)]. Most remarkably, the continuum of excitations moves towards higher energy with increasing magnetic field, and appears to be associated with the gapped excitation branch. Moreover, the degeneracy at the M point is completely lifted, with the energy of the excitation in the gapped branch increased to 0.35 meV [Figure 11(g)]. The roton-like dip in the gapped branch becomes hardly visible [Figure 10(c)]. Overall, the excitation intensity becomes much weaker at this field (note the $\times$ 2 factor in color scale).
Finally, the continuum of excitations disappears at $\mu_{0}H=0.75$ T, just below the quantum phase transition to the uud plateau phase at $\mu_{0}H_{c}\approx 0.8$ T, as shown in Figure 10(d). The excitations evolve into sharp magnons, resembling those observed in the plateau phase at $\mu_{0}H_{c}=1.5$ T, as previously reported [33]. The gapless branch is well described by LSWT (solid lines). It is worth noting that, LSWT predicts a lower critical field implying that the calculated magnon bands in Figure 10(d) are gapped. This reaffirms that the continuum of excitations in the supersolid state of K2Co(SeO3)2 is linked to the “superfluidity” of the transverse $xy$-moment in the Y-structure.
In contrast, the gapped branch exhibits minor deviations from the LSWT predictions [Figure 10(d)]. These discrepancies arise because LSWT is not fully accurate near the quantum critical point. Although the spin structure closely resembles the collinear uud configuration, quantum fluctuations significantly reduce the ordered moment from its saturated value [33]. Furthermore, the scattering appears diffusive: the energy width of the excitations slightly exceeds the resolution [Figure 11(h)], and additional weak scattering is observed at the K point near 0.5 meV, just below the gapped excitation branch [see Figure 10(d) and Figure 11(d)]. To further highlight the distinctions of the measured excitation spectra from LSWT, the dynamical spin structure factor (DSSF) calculated using LSWT are shown in Supplementary Figure 6 (left 4 panels).
II.4 Hamiltonian model